# Characterizing Gaussian quantum processes with Gaussian resources

### v1

@@ I Introduction
A variety of tasks in quantum information science can be implemented by continuous variable systems, of which quantum light is a leading platform Andersen et al. (2010); Braunstein and van Loock (2005). Such implementations necessitate the production, transformation, and detection of quantum light. Recently, several platforms have been developed for these tasks Raffaelli et al. (2018); Madsen et al. (2022). Due to manufacturing imperfections, a post-production characterization of these technologies is necessary to ensure that such devices function as intended Tan and Rohde (2019).
Characterizing continuous variable systems can be inherently challenging, as the Hilbert space corresponding to such systems is infinite-dimensional Weedbrook et al. (2012). However, a simplification is possible when we restrict our attention to Gaussian systems. These systems include linear optical systems as well as those with quadratic nonlinearities such as parametric amplifiers and down-converters, and squeezers Caves (1982); Braunstein (2005). Despite this restriction, Gaussian systems enable the execution of several important quantum information tasks including communication, cryptography, teleportation, and non-universal quantum computation such as Gaussian boson sampling Weedbrook et al. (2012).
@@
Several investigations have focused on characterizing linear optical systems by determining their $N\times N$ unitary transfer matrix. This was done using non-Gaussian probes in Refs. Peruzzo et al. (2011); Laing and O’Brien (2012); Dhand et al. (2016); Zhou et al. (2015), and with Gaussian probes in Refs. Rahimi-Keshari et al. (2013); Suess et al. (2020); Hoch et al. (2023). However, these schemes assumed that the unitary matrix is real-bordered, i.e. the elements in its first row and first column are real. This restriction was removed in Ref. Jacob et al. (2018), which considered not only linear optical systems but also those with quadratic non-linearities i.e. all Gaussian systems. However, non-classical probes were still used in this analysis.
Several other works have explored beyond linear optics by characterizing all Gaussian systems. In Ref. Wang et al. (2013), the Husimi Q-function corresponding to the Gaussian process was found using coherent probes. In contrast, coherent probes and photon number-resolving measurements were used to directly find the parameters of the Gaussian map in Ref. Kumar et al. (2020). Ref. Teo et al. (2021) uses coherent states and heterodyne measurements to characterize a single-mode Gaussian process.
Despite this progress, a method that directly characterizes multi-mode Gaussian processes with only Gaussian resources such as coherent probes and quadrature detection is still needed. In this paper, we develop such a method. Using this method, we directly obtain all elements of the symplectic matrix that completely describes the process.
This paper is structured as follows: In section II, we outline the mathematical conventions used in this work. Then in section III, we give details of our proposed setup. Section IV then shows the procedure employed in our method. In section V, we incorporate loss in devices and demonstrate that the scheme is resilient to uniform loss. The results of simulating our method are presented in section VI. We discuss the results in section VII, and conclude in section VIII.
@@ II Quantum Continuous Variable Systems
Consider a $N-$mode bosonic quantum system that has the corresponding mode annihilation and creation operators $\{\hat{a}_{i},\hat{a}_{i}^{\dagger}\}_{i=1}^{N}$. These operators satisfy the commutation relation $\left[\hat{a}_{i},\hat{a}_{j}^{\dagger}\right]=\delta_{ij}$. These operators can be used to define a set of quadrature operators for each mode as
@@
${\bf{\hat{r}}}\equiv\left(\hat{X}_{1},\dots,\hat{X}_{N},\hat{P}_{1},\dots,\hat{P}_{N}\right)^{T}.$ (2)
@@ II.1 Gaussian States
${\bf{r}}\equiv\left\langle{\bf{\hat{r}}}\right\rangle=\mathrm{Tr}\left({\bf{\hat{r}}}\hat{\rho}\right).$ (3)
@@
[…]
@@ II.2 Gaussian Processes
where $S$ is a $2N\times 2N$ invertible symplectic matrix of real elements with determinant $+1$ Weedbrook et al. (2012); Adesso et al. (2014); Serafini (2017). Since $S$ is symplectic, we have
@@
If the Gaussian quantum process corresponds to passive transformations of the mode operators, as in linear optics, the $2N\times 2N$ symplectic matrix $S$ is related to a $N\times N$ unitary matrix $U$ that transforms the mode operators of the linear optical system Valson Jacob (2020). It is shown in Appendix A that $U$ is related to $S$ as
@@ II.3 Gaussian measurements
A quantum measurement is defined to be a gaussian measurement if its application to a gaussian state yields a gaussian-distributed outcome Weedbrook et al. (2012). Two common gaussian measurements are the homodyne and the heterodyne. Both of these are single-mode measurements that can be used to reconstruct the first moment (i.e. the mean vector) of a quantum state.
Homodyne detection is modeled as a projection onto the quadrature basis Weedbrook et al. (2012). As such, it can measure any given quadrature of a quantum state at a time. In contrast, heterodyne detection is modeled as a projection onto the coherent basis. Thus, it jointly measures complementary quadratures and is subject to the unavoidable Heisenberg-Robertson-Schrödinger uncertainty Teo et al. (2017).
@@ III Setup
Figure 1: A schematic for characterizing Gaussian devices. A coherent probe $|\alpha\rangle$ is input into the $j^{\text{th}}$ mode of a device $D$. After the probe evolves through the device, output quadratures are measured in each output mode.
As shown in Fig. 1, the setup consists of a Gaussian device $D$ with $N$ input and output modes. A well-characterized coherent probe is sent into the $j^{\text{th}}$ input mode of the device and vacuum states are input into every other mode. The coherent probe then transforms through the device following the map of Eq. (5). Quadratures are measured at the $i^{\text{th}}$ output mode, either by homodyne detection or by heterodyne detection. This process is repeated for all combinations of $i,j\in\{1,2,\ldots,N\}$.
@@ IV Characterization of Gaussian processes
[…]
@@
$S_{i,j}=\frac{\langle\hat{X}_{i}\rangle^{\text{out}}}{\sqrt{2}\alpha}$ (11)
@@
Finally, we determine all elements of $S$ by choosing all combinations of input modes $j$ and output modes $i$. Fig. 2 shows the measurement settings necessary to find all elements of $S$, requiring $4N^{2}$ measurement settings in total.
Figure 2: Figure shows the probes and measurement settings needed to find the elements of the matrix $S$. R denotes row number and C denotes column number. Four elements of the matrix can be determined for a given input and output mode by changing the phase of the probe as well as the quadrature measured.
IV.1 Simplification for passive transformations
@@
V Modeling Loss
In this section, we account for loss that is inevitable in any experiment. We consider a uniform (or balanced) loss model, where loss is the same in each mode. Such loss is modeled by a set of fictitious beam splitters of the same transmissivity $\eta\in(0,1]$ attached to each input mode of the device Gerry and Knight (2023). Vacuum is input into all of the ancillary modes. A schematic of this loss model is shown in Fig. 3.
@@ V Modeling Loss
It was shown in Ref. Oszmaniec and Brod (2018) that uniform loss commutes with passive linear optical transformations. As shown in Appendix B, uniform loss commutes with the Gaussian map of Eq. (5) also. Therefore, this loss model is agnostic to the position of loss in any device with uniform loss. This allows us to transfer any uniform loss within the device to its input. Thus, the effect of uniform loss is to attenuate the input coherent probe $|\alpha\rangle_{j}$ to $|\sqrt{\eta}\,\alpha\rangle_{j}$. Consequently, Eq. (9) is modified to include the loss as
${\bf{r}}^{\text{out}}=\sqrt{\eta}\langle\hat{X}_{j}\rangle^{\text{in}}\left(\begin{array}[]{c}S_{1,j}\\ S_{2,j}\\ \vdots\\ S_{2N,j}\end{array}\right)+\sqrt{\eta}\langle\hat{P}_{j}\rangle^{\text{in}}\left(\begin{array}[]{c}S_{1,N+j}\\ S_{2,N+j}\\ \vdots\\ S_{2N,N+j}\end{array}\right)$ (12)
@@
Thus, $\eta$ is directly determined from experimental data, thereby requiring no additional measurements. Once determined, $\eta$ is used to obtain $S$ from the experimentally obtained matrix $\tilde{S}$ as $S=\frac{1}{\sqrt{\eta}}\tilde{S}$. This procedure is summarized in Algorithm 1.
@@ VI Simulation
To test the validity of our method, we simulate it on Strawberry Fields, an open-source Python library for quantum continuous variable systems Killoran et al. (2019); Bromley et al. (2020). The corresponding Python code used for the simulation is available on GitHub at Grove and Valson Jacob (2025).
We randomly generated symplectic matrices $S^{\text{rand}}$ for $N-$mode devices where $N$ varied from 2 to 20. For each case, we implemented the corresponding Gaussian transformation and uniform loss. We then reconstructed those devices using either homodyne or heterodyne detection to obtain $S^{\text{recon}}$. The accuracy of reconstruction was quantified by a scaled Frobenius (Hilbert–Schmidt) norm of the difference between the randomly generated matrix and the reconstructed matrix. This norm is defined as
@@
$=\frac{1}{N}\sqrt{\sum_{i,j}^{N}\left|S^{\text{rand}}_{i,j}-S^{\text{recon}}_{i,j}\right|^{2}},$ (14)
where $N$ is the number of modes. A lower value of $F$ indicates a more accurate reconstruction. We repeated this process 500 times and calculated the average of this norm for each choice of $N$.
In our simulation, we used coherent probes of amplitude $\alpha=1000$. The transmissivity $\eta$ was set at either 0% or at 50%. We ensured that the same number of coherent probes were used in both homodyne and heterodyne reconstruction. This was necessary because homodyne reconstruction requires two probes to measure both quadratures while heterodyne only requires one.
The results of the simulation are shown in Fig. 4. We note that the value of the scaled Frobenius norm is independent of the number of modes of the device. This indicates that our method is readily scalable with the size of the reconstructed device. We also notice that heterodyne reconstruction outperforms homodyne reconstruction. This is supported by two observations. First, for each value of the loss parameter, heterodyne reconstruction yields a lower value of the scaled Frobenius norm than that of homodyne reconstruction. Second, the uncertainties in the homodyne data are larger than those in heterodyne data.
Figure 4: Average scaled Frobenius norm of the difference between the randomly generated symplectic matrix and its reconstruction. Reconstruction was performed using both heterodyne and homodyne measurements. The loss parameter was set either at 0% or at 50%. Heterodyne reconstruction is observed to outperform homodyne reconstruction.
Figure 5: Average scaled Frobenius norm of the difference between the randomly generated unitary matrix and its reconstruction. Reconstruction was performed using both heterodyne and homodyne measurements for lossless and lossy linear devices. The norm is observed to be similar for both homodyne and heterodyne reconstruction.
@@
VII Discussion
The results of the simulation in Sec. VI demonstrate the superiority of heterodyne measurements for reconstructing Gaussian processes but not for linear processes. We now consider why this might be the case.
We suggest that the superiority of heterodyne reconstruction is explained by the nature of the quantum state on which quadrature measurements are performed. In the context of quantum state reconstruction, it was shown in Ref. Teo et al. (2017) that both heterodyne detection and homodyne detection perform equally well in reconstructing the first moments of any minimum-uncertainty state. However, if the state is not a minimum-uncertainty state, heterodyne reconstruction performs better. This result is directly relevant to our method, as our reconstruction is dependent on measuring the first moment of the output state of the unknown process.

### v2

@@ I Introduction
A variety of tasks in quantum information science can be implemented by continuous variable systems, of which quantum light is a leading platform [1, 2]. Such implementations necessitate the production, transformation, and detection of quantum light. Recently, several platforms have been developed for these tasks [3, 4]. Due to manufacturing imperfections, a post-production characterization of these technologies is necessary to ensure that such devices function as intended [5].
Characterizing continuous variable systems can be inherently challenging, as the Hilbert space corresponding to such systems is infinite-dimensional [6]. However, a simplification is possible when we restrict our attention to Gaussian systems. These systems include linear optical systems as well as those with quadratic nonlinearities such as parametric amplifiers and down-converters, and squeezers [7, 8]. When used in conjunction with non-Gaussian systems, Gaussian systems enable the execution of several important quantum information tasks including communication, cryptography, teleportation, and non-universal quantum computation such as Gaussian boson sampling [6].
@@
Several investigations have focused on characterizing linear optical systems by determining their $N\times N$ unitary transfer matrix. This was done using non-Gaussian probes in Refs. [9, 10, 11, 12] and with Gaussian probes in Refs. [13, 14, 15]. However, these schemes assumed that the unitary matrix is real bordered, i.e. the elements in its first row and first column are real. This restriction was removed in Ref. [16], which considered not only linear optical systems but also those with quadratic nonlinearities i.e. all Gaussian systems. However, nonclassical probes were still used in this analysis.
Several other works have explored beyond linear optics by characterizing all Gaussian systems. In Ref. [17] the Husimi $\it{Q}$ function corresponding to a Gaussian process was found using coherent probes. In contrast, coherent probes and photon number-resolving measurements were used to directly find the parameters of the Gaussian map in Ref. [18]. Reference  [19] uses coherent states and heterodyne measurements to characterize a single-mode Gaussian process.
Despite this progress, a method that directly characterizes multimode Gaussian processes with only Gaussian resources is still needed. In this paper, we develop such a method, using no nonclassical resources, and employing only coherent probes and quadrature detection. Using this method, we directly obtain all elements of the symplectic matrix that completely describes the process.
This paper is structured as follows: In Sec. II we outline the mathematical conventions used in this work. Then in Sec. III we give details of our proposed setup. Section IV then shows the procedure employed in our method. In section V, we consider experimental limitations that can constrain implementations of the proposed method. The results of simulating our method are presented in Sec. VI. We discuss salient results in Sec. VII, and conclude in Sec. VIII.
@@ II Quantum Continuous Variable Systems
Consider an $N$-mode bosonic quantum system that has the corresponding mode annihilation and creation operators $\{\hat{a}_{i},\hat{a}_{i}^{\dagger}\}_{i=1}^{N}$. These operators satisfy the commutation relation $\left[\hat{a}_{i},\hat{a}_{j}^{\dagger}\right]=\delta_{ij}$. These operators can be used to define a set of quadrature operators for each mode as
@@
${\bf{\hat{r}}}\equiv\left(\hat{X}_{1},\dots,\hat{X}_{N},\hat{P}_{1},\dots,\hat{P}_{N}\right)^{\mathrm{T}}.$ (2)
@@ II.1 Gaussian states
[…]
@@
where $\{.,.\}$ represents the anticommutator. Examples of Gaussian states are the vacuum state, coherent states, thermal states, and squeezed states [20].
@@ II.2 Gaussian processes
where $S$ is a $2N\times 2N$ invertible symplectic matrix of real elements with determinant +1+1 [6, 21, 22]. Since $S$ is symplectic, we have
@@
If the Gaussian quantum process corresponds to passive transformations of the mode operators, as in linear optics, the $2N\times 2N$ symplectic matrix $S$ is related to a $N\times N$ unitary matrix $U$ that transforms the mode operators of the linear optical system [23]. It is shown in Appendix A that $U$ is related to $S$ as
@@ II.3 Gaussian measurements
A quantum measurement is defined to be a Gaussian measurement if its application to a Gaussian state yields a Gaussian-distributed outcome [6]. Two common Gaussian measurements are the homodyne and the heterodyne. Both of these are single-mode measurements that can be used to reconstruct the first moment (i.e. the mean vector) of a quantum state.
Homodyne detection is modeled as a projection onto the quadrature basis [6]. As such, it can measure any given quadrature of a quantum state at a time. In contrast, heterodyne detection is modeled as a projection onto the coherent basis. Thus, it jointly measures complementary quadratures, and is subject to additional Arthurs-Kelly type noise that is inherent to such joint measurements [24].
@@ III Setup
Figure 1: A schematic for characterizing Gaussian devices. A coherent probe $|\alpha\rangle$ is input into the $\it{j}{\text{th}}$ mode of a device $D$. After the probe evolves through the device, output quadratures are measured in each output mode.
As shown in Fig. 1, the setup consists of a Gaussian device $D$ with $N$ input and output modes. A well-characterized coherent probe is sent into the $\it{j}{\text{th}}$ input mode of the device and vacuum states are input into every other mode. The coherent probe then transforms through the device following the map of Eq. (5). Quadratures are measured at the $\it{i}{\text{th}}$ output mode, either by homodyne detection or by heterodyne detection. This process is repeated for all combinations of $i,j\in\{1,2,\ldots,N\}$.
@@ IV Characterization of Gaussian processes
${\bf{r}}^{\text{in}}=\left(0,\;\cdots,\;\langle\hat{X}_{j}\rangle^{\text{in}},\;0,\;\cdots,\;\langle\hat{P}_{j}\rangle^{\text{in}},\;\cdots,\;0\right)^{\text{T}}.$ (8)
@@
$S_{i,j}=\frac{\langle\hat{X}_{i}\rangle^{\text{out}}}{\sqrt{2}\alpha}.$ (11)
@@
Finally, we determine all elements of $S$ by choosing all combinations of input modes $j$ and output modes $i$. Figure  2 shows the measurement settings necessary to find all elements of $S$, requiring $4N^{2}$ measurement settings in total.
Figure 2: The probes and measurement settings needed to find the elements of the matrix $S$. R denotes row number and C denotes column number. Four elements of the matrix can be determined for a given input and output mode by choosing the phase of the probe as well as the quadrature measured.
Simplification for passive transformations
@@
V Modeling experimental constraints
In this section, we account for two realistic limitations that can constrain implementations of the proposed method: optical loss and phase-modulation errors.
V.1 Optical loss
Realistically, optical loss is inevitable to any experiment. Such loss may occur while light is input into the components of the device (input loss) as well as while light propagates through the device (propagation loss) [25]. To model these, we consider a uniform (or balanced) loss model, where loss is the same in each mode. Such loss is modeled by a set of fictitious beam splitters of the same transmissivity $\eta\in(0,1]$ attached to each input mode of the device [26]. Vacuum is input into all of the ancillary modes. A schematic of this loss model is shown in Fig. 3.
@@ V.1 Optical loss
It was shown in Ref. [27] that uniform loss commutes with passive linear optical transformations. As shown in Appendix B, uniform loss commutes with the Gaussian map of Eq. (5) also. Therefore, this loss model is agnostic to the position of loss in any device with uniform loss. This allows us to transfer any uniform loss within the device to its input. Thus, the effect of uniform loss is to attenuate the input coherent probe $|\alpha\rangle_{j}$ to $|\sqrt{\eta}\,\alpha\rangle_{j}$. Consequently, Eq. (9) is modified to include the loss as
${\bf{r}}^{\text{out}}=\sqrt{\eta}\langle\hat{X}_{j}\rangle^{\text{in}}\left(\begin{array}[]{c}S_{1,j}\\ S_{2,j}\\ \vdots\\ S_{2N,j}\end{array}\right)+\sqrt{\eta}\langle\hat{P}_{j}\rangle^{\text{in}}\left(\begin{array}[]{c}S_{1,N+j}\\ S_{2,N+j}\\ \vdots\\ S_{2N,N+j}\end{array}\right).$ (12)
@@
Thus, $\eta$ is directly determined from experimental data, thereby requiring no additional measurements. Once determined, $\eta$ is used to recover $S$ from the experimentally obtained matrix $\tilde{S}$ as $S=\frac{1}{\sqrt{\eta}}\tilde{S}$. This procedure is summarized in Algorithm 1.
@@
V.2 Phase-modulation errors
As the method proposed requires the precise modulation of the phases of the input probes, we investigate the effect of phase-modulation errors. To this end, without loss of generality, we consider the reconstruction of a symplectic matrix element $S_{i,j}$ where $i,j\in\{1,2,\ldots,N\}$. The probe used is labeled $|\alpha=re^{i\phi}\rangle$, where $r\in\mathbb{R}$ and $\phi$ models the phase-modulation error. From Eq. (9), the measured output quadrature is
$\langle\hat{X}_{i}\rangle^{\text{out}}$ $=\langle\hat{X}_{j}\rangle^{\text{in}}S_{i,j}+\langle\hat{P}_{j}\rangle^{\text{in}}S_{i,N+j}$
$=\sqrt{2}r\cos{\phi}\,S_{i,j}+\sqrt{2}r\sin{\phi}\,S_{i,N+j}$
$=\sqrt{2}rS_{i,j}+\sqrt{2}r\phi S_{i,N+j}+O(\phi^{2}).$ (14)
This relation enables us to find $S_{i,j}$ as
$S_{i,j}=\frac{\langle\hat{X}_{i}\rangle^{\text{out}}}{\sqrt{2}r}-\phi S_{i,N+j}+O(\phi^{2}).$ (15)
Comparing Eq. (15) with Eq. (11), we note that phase-modulation error $\phi$ introduces a first-order error in the reconstructed matrix element. If we further assume $\phi$ to be randomly distributed with a mean value of zero, then averaging our results over many trials suppresses the reconstruction error to second order in $\phi$. Thus the reconstruction is robust to small phase-modulation errors.
@@ VI Simulation
To test the validity of our method, we simulate it on Strawberry Fields, an open-source Python library for quantum continuous variable systems [28, 29]. The corresponding Python code used for the simulation is available on GitHub [30]. For each simulation, the accuracy of reconstruction was quantified by a scaled Frobenius (Hilbert–Schmidt) norm of the difference between a randomly generated symplectic matrix $S^{\text{rand}}$ and its reconstruction $S^{\text{recon}}$. This norm is defined as
@@
$=\frac{1}{N}\sqrt{\sum_{i,j}^{2N}\left|S^{\text{rand}}_{i,j}-S^{\text{recon}}_{i,j}\right|^{2}},$ (16)
where $N$ is the number of modes. A lower value of $F$ indicates a more accurate reconstruction.
VI.1 Scalability with the number of modes
To demonstrate the scalability of the reconstruction with the number of modes, we randomly generated symplectic matrices $S^{\text{rand}}$ for $N$-mode devices where $N$ varied from 2 to 20. For each case, we implemented the corresponding Gaussian transformation as well as uniform loss. We then reconstructed those devices using either homodyne or heterodyne detection to obtain $S^{\text{recon}}$. Finally, we calculated the accuracy of reconstruction $F$.
To eliminate any bias in our initial choice of $S^{\text{rand}}$, we repeated this process 500 times and calculated the average value of $F$ for each choice of $N$. In our simulation, we used coherent probes of amplitude $\alpha=1000$. The transmissivity $\eta$ was set at either 0% or at 50%. We ensured that the same number of coherent probes were used in both homodyne and heterodyne reconstruction. This was necessary because homodyne reconstruction requires two probes to measure both quadratures while heterodyne only requires one.
The results of the simulation are shown in Fig. 4. We note that $F$ tends to be independent of the number of modes of the device. This indicates that our method is readily scalable with the size of the reconstructed device. We also notice that heterodyne reconstruction outperforms homodyne reconstruction. This is supported by two observations. First, for each value of the loss parameter, heterodyne reconstruction yields a lower value of the scaled Frobenius norm than that of homodyne reconstruction. Second, the uncertainties in the homodyne data are larger than those in heterodyne data.
Figure 4: Accuracy of reconstruction as a function of the number of modes. Reconstruction was performed using both heterodyne and homodyne measurements. The loss parameter was set either at 0% or at 50%. Heterodyne reconstruction is observed to outperform homodyne reconstruction.
Figure 5: Accuracy of reconstruction as a function of the number of modes. Reconstruction was performed using both heterodyne and homodyne measurements for lossless and lossy linear devices. The accuracy is observed to be similar for both homodyne and heterodyne reconstruction.
@@ VI.1 Scalability with the number of modes
VI.2 Scalability with the coherent probe intensity
In the proposed method, we obtain the elements of the unknown symplectic matrix as a ratio of the output and input quadratures, as seen in Eq. (11). Therefore, the input probe intensity will affect the accuracy with which we determine the unknown symplectic matrix elements. A larger probe intensity shall increase the accuracy of reconstructing individual matrix elements, as can be noted from Eq. (11). This limit in the accuracy of reconstruction follows the standard quantum limit as we use only classical resources in the reconstruction [31].
As experimental constraints may restrict the intensity of coherent probes used, one may wonder if the reconstruction method is accurate only with sufficiently strong coherent probes. We answer this question by demonstrating that the reconstruction method works for arbitrarily weak coherent probes. While using weak coherent probes, we repeat the method over multiple trials and average over the recovered matrix element. This repetition alone can reduce the standard error in measurements that scales inversely with $\sqrt{n}$ where $n$ is the number of trials.
To demonstrate this numerically, we randomly generated the symplectic matrix corresponding to a five-mode Gaussian device. We reconstructed it using heterodyne measurements for a range of coherent probe intensities. The reconstruction was performed using a single trial as well as with a multiple number of trials. The accuracy of reconstruction was calculated for each case, and plotted against the coherent probe intensity in Fig. 6.
Figure 6: Accuracy of reconstruction as a function of the the intensity of the coherent probes used. A 5-mode Gaussian device was randomly generated and then reconstructed using heterodyne measurements. For each coherent probe intensity, reconstruction was performed using a single trial as well as with a variable number of trials $n$. The accuracy of reconstruction is the same in cases where $n|\alpha|^{2}$ is the same.
In Fig. 6, we note that the accuracy of reconstruction is the same for multiple coherent probe intensities provided that reconstruction with weaker probes was repeated multiple times. This is supported by the observation that the scaled Frobenius norm is the same as long as $n|\alpha|^{2}$ (or $\sqrt{n}|\alpha|$) remains the same. Since $n|\alpha|^{2}$ is a measure of the average number of photons used to reconstruct any given matrix element, the accuracy of reconstruction is limited by the energy of the probes alone. Therefore one may use weaker coherent probes to obtain the same reconstruction accuracy as long as it is used for a larger number of trials such that $n|\alpha|^{2}$ remains fixed.
VII Discussions
In this section we first comment on two salient observations: the superiority of heterodyne measurements in reconstructing Gaussian devices as seen in Sec.VI.1 and the utility of the reconstruction method beyond the Gaussian regime.
VII.1 Superiority of heterodyne measurements
The results of the simulation in Sec. VI.1 demonstrate the superiority of heterodyne measurements for reconstructing Gaussian processes but not for linear processes. We now consider why this might be the case.
We suggest that the superiority of heterodyne reconstruction is explained by the nature of the quantum state on which quadrature measurements are performed. In the context of quantum state reconstruction, it was shown in Ref. [24] that both heterodyne detection and homodyne detection perform equally well in reconstructing the first moments of any minimum-uncertainty state. However, if the state is not a minimum-uncertainty state, heterodyne reconstruction performs better. This result is directly relevant to our method, as our reconstruction is dependent on measuring the first moment of the output state of the unknown process.